Not sure which topic would fit this best... but if someone could show me how to tackle this question:
The sum of two numbers is 30 and their product is a maximum. Find the two numbers.

Answer by Alan3354(69443)

Not sure which topic would fit this best... but if someone could show me how to tackle this question:
The sum of two numbers is 30 and their product is a maximum. Find the two numbers.
------------
It's always the square, in this case 15 and 15.
15*15 = 225
14*16 = 224
13*17 = 221
See the pattern?
------------
A rigorous solution:
x*(30-x) = 0
x^2 - 30x = 0
The vertex (maximum) is at x = -b/2a
= 30/2 = 15
----------
Using calculus:
f(x) = x*(30-x)
f(x) = x^2 - 30x
f'(x) = 2x-30 = 0
2x-30 = 0
x = 15